# RISCORE: Enhancing In-Context Riddle Solving in Language Models through Context-Reconstructed Example Augmentation

### v3

@@ 4.2 Baselines
In our study, we used various prompting techniques as baselines, including zero-shot (ZS), few-shot (FS), and Chain of Thought (CoT) methods. A detailed analysis of each technique can be found in the App. E.
Chain-of-thought (CoT_ZS)
In this experiment, we prompt the model to answer the multiple-choice riddle in a step-by-step manner, following the method proposed in Wei et al. (2023), without including any examples. This approach allows us to assess the models’ riddle-solving abilities in a zero-shot setting.
@@
To assess the influence of example-based learning, we incorporated few-shot prompting with varying numbers of examples: 2, 4, and 8 shots. We limited our tests to 8 shots since it has been proven that reasoning abilities of LLMs are deteriorating with longer inputs Levy et al. (2024). Two distinct strategies were used for selecting these examples: random selection, referred to as Rand, and selecting riddles with minimal semantic distance from the test riddle using a language model Zhang et al. (2024) [^2] https://huggingface.co/Alibaba-NLP/gte-large-en-v1.5, referred to as Sim.
Few-shot with CoT Explanations (CoT_FS)
A straightforward way to aid the model in understanding an exemplar is by providing a supplementary explanation. In this experiment, we explored this approach by offering explanations alongside the few-shot exemplars. This allowed us to evaluate whether these crafted explanations could improve the models’ ability to comprehend and follow the required reasoning process. The explanations were generated using a semi-automated process. First, for each exemplar in the FS setting, we used ChatGPT [^3] Specifically, we used version gpt-3.5-turbo-0125. to generate an explanation for the riddle based on the correct answer, following the method proposed in Wei et al. (2023). These explanations were then manually curated to ensure alignment with human interpretations of the reasoning process required for the riddle. Due to the labor-intensive process of annotating the explanations, we conducted this experiment using only a randomly selected sampling strategy. Additionally, the results indicate that the explanations do not significantly aid the models in solving riddles (see Section 5 for further details).
@@ 4.3 RISCORE
RISCORE is applied as follows. For each example used in the few-shot (FS) setting, the corresponding contextually reconstructed riddle is appended to the input. For a fair comparison, the number of shots refers to the total number of examples included in the prompt, regardless of whether they are new riddles or context reconstructed examples. Thus, for example RISCORE with 4 shots has the same 2 examples used in 2-shots FS Sim technique augmented with the context reconstruction of them, meaning that keeps only half of the examples drawn from the original dataset and the remainder consisting of reconstructed examples. The experiments were conducted with 2, 4, and 8-shot settings, meaning that 1, 2, and 4 exemplars, respectively, were drawn from the dataset, with the rest being their context reconstructions.
We primarily utilized the Llama3-8B and Llama3-70B model to generate both Question-Answer pairs and distractors. The Question-Answer pairs were produced in both zero-shot and few-shot settings.
Lastly, in our experiments, we differentiate between RISCOREm, which utilizes manually created reconstructions (when available), and RISCORE, which employs the fully automated method for generating reconstructions. The prompt structure for the system-user interaction in this method is the same as in the FS method outlined in App. E. The sole distinction is in the examples provided.
@@
Table 1: Model’s performance for Brain Teaser using baselines and RISCOREm prompting. The best overall results are highlighted in bold. More detailed results can be found in Tables 6 and 7.
@@
Llama3-70B zeroshot for QA & Llama3-8B distractors
@@
Llama3-70B fewshot for QA & Llama3-8B distractors
@@
Llama3-70B fewshot for QA & Llama3-70B distractors
@@
Table 2: Model’s performance for Brain Teaser using RISCORE prompting. Similarity-based selection was employed for choosing all the exemplars. Results that surpass the FS method with semantically similar examples, using the same number of shots, are underlined. More detailed results can be found in Tables 7 and 8.
@@ 4.4 Models
To this end, we test various models of different scales. Specifically, we examine Llama3 with 8 and 70 billion parameters respectively AI@Meta (2024), Mistral-7B and Mistral-8x7B Jiang et al. (2023a), as well as Qwen2-7B Yang et al. (2024). This diverse set of models allows us to investigate the impact of contextually reconstructed examples on models with varying sizes and parameterizations. Further details regarding the model selection process and technical specifications can be found in App. F.
@@ 5 Results
In this section, we present and analyze the results of our experiments. Additional results, utilizing different metrics and hyperparameter settings, are available in App. G.
5.1 BrainTeaser
Table 2 presents the performance outcomes of various prompting techniques applied to the BrainTeaser dataset. Notably, the CoT_FS method underperforms compared to few-shot techniques, even when supplied with examples that include manually crafted explanations. This trend holds regardless of the model size utilized. Conversely, there is a significant improvement when the few-shot approach incorporates the provided “context reconstructed” riddles alongside the original riddles. This suggests that context reconstructed riddles effectively guide the model in tracing the appropriate reasoning path. Remarkably, the benefit of using these reconstructed contexts is more pronounced in smaller models compared to larger ones of the same architecture. For instance, the improvement in the Llama3-8B model is substantial: while the 4-shot technique using semantic similarity achieves a score of 0.717, augmenting this approach with four context reconstructed examples (totaling eight shots) enhances performance to 0.758—an improvement of approximately 5.7%. For the larger model, this increase is from 0.792 to 0.808, leading to an improvement of approximately 2.02%. The same pattern is more evident to the Mistral-7B, when for 4-shot we notice an increase of 9.2% while for Mixtral-8x7B we have an increase of 2.5%. A similar trend can be observed in Qwen2-7B, where the 4-shot setting with semantically similar examples performs approximately 3.4% lower than the augmented 8-shot exemplars, which include the same examples along with their context reconstructions, resulting in a total of 8 examples.
At this point, it is worth noting that the performance improvement does not originate solely from the increased number of shots but also from the quality of the shots added. This is evident from the fact that the best results on the dataset are achieved when we include examples with their context reconstruction riddles. Furthermore, by comparing performance while keeping the number of shots constant, we observe that cases using the reconstructed riddles perform better, especially when the number of shots is more than two. In the latter case, the addition of different examples seems more helpful to the model in properly understanding the task, since including only a single example with its contextual reconstruct makes the prompt overly specific.
@@ 5.1 BrainTeaser
Table 2 presents the performance of RISCORE using the automated method for generating context reconstructed examples. As expected, the using the automated method yields a smaller improvement in model performance compared to the manually curated examples provided in the BrainTeaser dataset. However, the performance of all models was consistently improved by the addition of context reconstructed riddles, even when using the fully automated algorithm for their generation. For example, for the Llama3-8B model, the performance using 4 shots from the dataset is 0.717, and by using the same shots along with their reconstructions generated by the automated method (8 shots in total), the performance increases to 0.742. The same trend is evident across all relatively smaller models, where augmenting the 4-shot examples selected based on semantic similarity with their automated context pairs yields notable performance improvements. Specifically, in the case of Qwen2-7B, the 4-shot setting augmented with corresponding context exemplars results in a 2.5% performance improvement. In Mistral-7B, the improvement is even more pronounced, with performance increasing by 10% from a baseline of 0.458 in the best-case scenario. Importantly, Table 2 shows that in most cases (underlined), it is even better to use the automated method to generate additional examples compared to adding the same number of examples from the dataset. For instance, the performance of Llama3-70B with 8 shots from the dataset is 0.783, whereas using only 4 examples from the dataset and automatically creating the remaining 4 using the context reconstruction method results in a performance of 0.808.
@@
Table 3: Model’s Performance for RiddleSense using baseline techniques. The best results overall are in bold. More detailed results can be found in Table 9.
@@
Table 4: Model’s performance for RiddleSense using RISCORE prompting. Similarity-based selection was employed for choosing all the exemplars. Results that surpass the FS method with semantically similar examples, using the same number of shots, are underlined. More detailed results can be found in Table 10.
@@ 5.2 RiddleSense
Table 4 presents the results of the baseline techniques using various models in the RiddleSense dataset. Once again, the results confirm that the few-shot technique, utilizing contextually similar exemplars for in-context learning, consistently delivers the best performance across all tested models.
In Table 4, we present the RiddleSense results using the proposed method for context reconstruction of each input. A clear trend emerges when comparing a simple 8-shot exemplar selection based on semantic similarity with the 8-shot setting of our method, where the top 4 semantically similar examples are augmented with our generated contextual pairs. Notably, the results show that our method consistently outperforms the standard 8-shot exemplar approach, demonstrating a significant improvement in model performance across various instances. An example of this trend is observed with the Llama3-8B model, where our method scores 0.708, approximately 2% higher than the few-shot setting based on semantic similarity, which achieves a score of 0.681. The same pattern is evident when comparing the two 4-shot settings. We achieve similar or marginally better accuracy using a total of just 4 examples—two original and two generated contextual reconstructions. This underlines the effectiveness of integrating contextually reconstructed pairs in enhancing model accuracy. We do not achieve a significant performance boost; however, we attain similar or marginally better results while relying on less grounded knowledge. This demonstrates the efficiency of our method, as it maintains comparable performance with fewer, yet strategically selected, exemplars.
@@ 5.3 Quality of context reconstructed riddles
To generate context reconstructed riddles, we used Llama3 models with 8 billion and 70 billion parameters in both few-shot (FS) and zero-shot (ZS) settings. We found that the Llama3-8B model struggled to produce high-quality Question-Answer pairs for the BrainTeaser dataset and was therefore not used in RISCORE. This difficulty likely arises from the BrainTeaser dataset’s demand for lateral thinking, which is particularly challenging for the smaller model. The Question-Answer pairs are essential, and if their quality is insufficient—as observed with Llama3-8B on the BrainTeaser dataset—high-quality distractors alone cannot compensate for this deficiency. However, for vertical thinking riddles, the smaller model effectively generates riddles even in the ZS setting, which, when used in the few-shot (FS) setting, can lead to increased performance compared to real examples drawn from the dataset.
@@ A.2 Prompts
First, we will present the zero-shot prompt configuration used for generating the contextually reconstructed Question-Answer pairs.
@@
  Next, with minor modifications, we introduce the few-shot prompt configuration that follows the same logic but incorporates in-context examples to guide the model’s responses.
@@ B.2 Riddlsense
The first pipeline involves a more granular approach by splitting the contextual reconstructed question into subphrases based on punctuation or conjunctions. If this yields fewer than three distinct subphrases, we further split the sentence at the position of the word “and” to create additional segments. We also detect the presence of interrogative words, and if the question is purely descriptive without any direct question, we append the phrase “What am I?” at the end. This is not chosen arbitrarily but follows the standard structure of many riddles in this dataset, where "What am I?" is the common question leading to single-word answers. Now that we have broken the riddle into sub-phrases and appended the appropriate question, we can prompt the model to generate a wrong answer for each sub-phrase concatenated with the question. This method ensures that the distractors align with different parts of the riddle’s setting. However, since some sub-phrases may contain key ideas central to the riddle, the model may still generate answers too similar to the correct one. To address this issue, we incorporate an additional intermediate step utilizing zero-shot classification. Specifically, we use the facebook/bart-large-mnli model 444facebook/bart-large-mnli Lewis et al. (2019) provided by Hugging Face. This model classifies the correct answer into one of eight general categories: ’food’, ’person’, ’object’, ’animal’, ’nature’, ’time’, ’place’, ’concept’. These categories are chosen to be mutually exclusive to avoid overlap. In our approach, we first predict the category of the correct answer using this model. We then leverage this classification to guide the generation of distractors. For each sub-phrase concatenated with the question, we provide the LLM with the two most similar categories (excluding the correct answer’s category). The model is prompted with a system-user instruction to produce a correct answer of the given setting in the given category. This method ensures that the generated distractors are not only incorrect but also contextually aligned with the riddle, thus maintaining the challenge for the model. After applying filtering to ensure the distractors are distinct and relevant, we have produced several distractors. The LLMs utilized in this pipeline include the previously mentioned Llama3-8B AI@Meta (2024) and its larger counterpart, Llama3-70B AI@Meta (2024).

### v4

@@ 4.2 Baselines
In our study, we use various prompting techniques as baselines, including zero-shot (ZS), few-shot (FS), and Chain of Thought (CoT) methods. A detailed analysis of each technique is provided in App. E.
Chain-of-thought (CoT ZS)
In this experiment, we prompt the model to answer the multiple-choice riddle in a step-by-step manner, following the method proposed in Wei et al. (2023), without demonstrating any examples. This approach allows us to assess the models’ riddle-solving abilities in a zero-shot setting.
@@
To assess the influence of example-based learning, we incorporate FS prompting with varying numbers of examples: 2, 4, and 8 shots. We limit our tests to 8 shots since it has been proven that reasoning abilities of LLMs are deteriorating with longer inputs Levy et al. (2024). Two distinct strategies were used for selecting these examples: random selection, referred to as Rand, and selecting riddles with minimal semantic distance from the test riddle using Zhang et al. (2024) [^3] https://huggingface.co/Alibaba-NLP/gte-large-en-v1.5 as the semantic similarity model, referred to as Sim.
Few-shot with CoT Explanations (CoT FS)
A straightforward way to aid the model in understanding an exemplar is by providing a supplementary explanation. In this experiment, we explore this approach by offering explanations alongside the few-shot exemplars. This allows us to evaluate whether these crafted explanations can improve the models’ ability to comprehend and follow the required reasoning process. The explanations are generated using a semi-automated process. First, for each exemplar in the FS setting, we use ChatGPT [^4] Specifically, we used version gpt-3.5-turbo-0125. to generate an explanation for the riddle based on the correct answer, following the method proposed in Wei et al. (2023). These explanations were then manually curated to ensure alignment with human interpretations of the reasoning process required for the riddle. Due to the labor-intensive process of annotating the explanations, this experiment is conducted using only the 2, 4, and 8-shot settings with a randomly selected sampling strategy.
@@ 4.3 RISCORE
RISCORE is applied as follows. For each example used in the FS setting, the corresponding contextually reconstructed riddle is appended to the input. For a fair comparison, the number of shots refers to the total number of examples included in the prompt, regardless of whether they are new riddles retrieved from the dataset or contextually reconstructed examples. Thus, for example RISCORE with 4 shots has the same 2 examples used in the 2-shot FS technique augmented with the 2 contextual reconstructed samples of them; ultimately, RISCORE keeps only half of the examples drawn from the original dataset, while the remainder consists of reconstructed examples. Therefore, the 2, 4, and 8-shot RISCORE settings correspond to 1, 2, and 4 exemplars drawn from the dataset respectively, with the rest of the demonstrations being their context reconstructions.
We primarily utilize Llama3-8B and Llama3-70B to generate both Question-Answer pairs and distractors. The Question-Answer pairs are produced in both ZS and FS settings. More details regarding these techniques can be found in App. E.
Lastly, in our experiments, we differentiate between RISCOREm, which utilizes manually created reconstructions (when available, as in existing BrainTeaset context reconstructions), and RISCORE, which employs our fully automated method for generating reconstructions. The prompt structure for the system-user interaction in this method is the same as in the FS method outlined in App. E. The sole distinction is in the examples provided.
@@
Table 1: Model performance for BrainTeaser using baselines and RISCOREm prompting. The best FS results are underlined, while best overall results per model are highlighted in bold. More detailed results can be found in Tables 6 and 7.
@@
Llama3-70B ZS for QA & Llama3-8B distractors
@@
Llama3-70B FS for QA & Llama3-8B distractors
@@
Llama3-70B FS for QA & Llama3-70B distractors
@@
Table 2: Model performance for BrainTeaser using RISCORE prompting. Similarity-based selection was employed for choosing all the exemplars. Results that surpass the FS method with semantically similar examples (FS Sim, check Table 2), using the same number of shots, are underlined. More detailed results can be found in Tables 7 and 8.
@@ 4.4 Models
To this end, we test various models of different scales on their reasoning abilities. Specifically, we examine Llama3 with 8 and 70 billion parameters respectively AI@Meta (2024), Mistral-7B and Mistral-8x7B Jiang et al. (2023a), as well as Qwen2-7B Yang et al. (2024). This diverse set of models allows us to investigate the impact of contextually reconstructed examples on models with varying sizes and parameterizations. By treating each model as a black-box (we merely prompt them and gather their responses), RISCORE adapts seamlessly to both open-source and proprietary models, demonstrating its versatility across a wide range of LLMs. Further details regarding the model selection process and technical specifications can be found in App. F.
@@ 5 Results
In this section, we present and analyze the results of our experiments. Additional results utilizing different metrics and hyperparameter settings are available in App. G.
5.1 BrainTeaser results
Table 2 presents the performance outcomes of various prompting techniques applied to BrainTeaser data. A key observation is the underperformance of the CoT FS method relative to FS techniques, even when examples are supplemented with manually crafted explanations. This underperformance persists regardless of the model size. As expected Dong et al. (2024), exemplar selection based on semantic similarity (FS Sim) consistently improves results when chosen over random selection (FS Rand), highlighting the value of semantic relevance to shot selection. This is a significant reference point for RISCORE, since a suitable initial exemplar selection is critical in properly guiding model reasoning when enhanced with contextually reconstructed examples.
Importantly, RISCOREm consistently outperforms FS Sim across all 4-shot and 8-shot configurations, as shown in Table 2. For instance, when comparing models with the same number of examples—such as 4-shot FS Sim versus 2+2-shot RISCOREm, or 8-shot FS Sim versus 4+4-shot RISCOREm—RISCOREm demonstrates superior performance. The advantages of RISCOREm are particularly evident in its ability to maintain robust reasoning.
To better understand this advantage, we can examine the ability of RISCOREm to mitigate noise from suboptimal shot selections, a notable issue in the FS Sim results. For example, with the Llama3-70B model, the 2-shot FS Sim achieves a performance score of 0.825. However, adding two more semantically similar examples reduces the score to 0.792. In contrast, the 2+2-shot RISCOREm achieves a score of 0.833—a 4% improvement over FS Sim with the same number of examples (vs 4-shot). This trend is consistent across other models. Similarly, with the smaller Mistral-7B model, the 2-shot FS Sim score is 0.517, but adding two more examples reduces it to 0.458. In contrast, the 2+2-shot RISCOREm, using contextually reconstructed examples, achieves a score of 0.567, outperforming FS Sim under identical shot conditions (vs 4-shot). These results indicate that RISCOREm effectively mitigates the noise introduced by suboptimal shot selections.
@@ 5.1 BrainTeaser results
The manually curated RISCOREm results serve as a benchmark, showcasing the potential upper bound of performance gains achievable with high-quality, handpicked riddles and distractors. Despite this, the automated RISCORE method consistently enhances performance by effectively leveraging reconstructed contexts, as shown in Table 2. While the improvements using the automated RISCORE are understandably smaller than those achieved with manually curated examples (check RISCOREm results of Table 2), they are nonetheless consistent and significant across all models tested. For example, Llama3-70B demonstrates significant improvement when transitioning from FS Sim with 8 semantically similar shots (0.783) to the RISCORE-augmented setting with 8 shots (0.808), which incorporates the original examples and their automated reconstructions. This trend is evident across all relatively smaller models, with noticeable performance gains. For Qwen2-7B, a 2-shot and 4-shot RISCORE setup results in a 2.5% improvement compared to the same-shot FS Sim setting, while for Mistral-7B, the improvement is even more pronounced—rising by up to 10% from a baseline of 0.458 under certain conditions in the 4-shot setting.
These results underscore the value of context reconstruction, particularly when the number of shots is limited, demonstrating that automated reconstruction can effectively complement semantically chosen examples. Interestingly, the results reveal that in most cases, RISCORE’s approach of combining N/2 semantically selected examples with their automated reconstructions outperforms the mere placement of the same number of examples retrieved from the dataset (indicated by the underlined results). However, there are instances where FS Sim outperforms RISCORE, primarily due to the selection constraints of RISCORE. In cases where the semantic similarity algorithm identifies an optimal example ranked lower than N/2 in a full FS setup, RISCORE may miss these shots due to its focus on N/2 examples plus reconstructions. This limitation indicates that RISCORE’s performance can be constrained by the quality of initial semantic similarity rankings and the resulting shot selection.
@@
Table 3: Model Performance for RiddleSense using baseline techniques. The best results overall are in bold. More detailed results can be found in Table 9. Note that no RISCOREm numbers are reported, since RiddleSense does not contain any ground truth reconstructed riddles.
@@
Table 4: Model performance for RiddleSense using RISCORE prompting. Similarity-based selection was employed for choosing all the exemplars. Results that surpass the FS method with semantically similar examples, using the same number of shots, are underlined. More detailed results can be found in Table 10.
@@ 5.2 RiddleSense
Table 4 presents the results of the baseline techniques using various models in the RiddleSense dataset. Once again, the results confirm that the few-shot technique, utilizing semantically similar exemplars for in-context learning, consistently delivers the best performance across all tested models.
In Table 4, we present the RiddleSense results using the proposed method for context reconstruction of each input. A clear trend emerges when comparing a simple 8-shot exemplar selection based on semantic similarity with the 8-shot setting of our method, where the top 4 semantically similar examples are augmented with our generated contextual pairs. Notably, the results show that our method consistently outperforms the standard 8-shot exemplar approach, demonstrating a significant improvement in model performance across various instances. An example of this trend is observed with the Llama3-8B model, where our method scores 0.708, approximately 2% higher than the few-shot setting based on semantic similarity, which achieves a score of 0.681. In the case of Mistral-8x7B, RISCORE avhieves a score of 0.700, a 2.5% gain over FS Sim’s 0.675, further demonstrating its consistent advantage. The same pattern is evident when comparing the two 4-shot settings. We achieve similar or marginally better accuracy using a total of just 4 examples—two original and two generated contextual reconstructions. This underlines the effectiveness of integrating contextually reconstructed pairs in enhancing model accuracy. We do not achieve a significant performance boost; however, we attain similar or marginally better results while relying on less grounded knowledge. This demonstrates the efficiency of our method, as it maintains comparable performance with fewer, yet strategically selected, exemplars, shifting the focus towards reasoning relevance of exemplars rather than quantity.
@@ 5.3 Quality of context reconstructed riddles
To generate contextually reconstructed riddles, we use Llama3 models with 8 billion and 70 billion parameters in both FS and ZS settings. We find that the Llama3-8B model struggles to produce high-quality Question-Answer pairs for the BrainTeaser dataset and is therefore not suitable for RISCORE. This difficulty likely arises from the BrainTeaser dataset’s demand for lateral thinking, which is particularly challenging for smaller reasoners. The Question-Answer pairs are essential, and if their quality is insufficient—as observed when leveraging Llama3-8B on BrainTeaser—high-quality distractors cannot compensate for this deficiency on their own. To address this issue, our preprocessing and filtering process ensures that only high-quality contextual examples are retained, maintaining the effectiveness of the method without being compromised by low-quality generations. However, for vertical thinking riddles, the smaller model effectively generates riddles even in the ZS setting, which, when used in the FS setting, can lead to increased performance compared to real examples drawn from the dataset.
@@ 6 Conclusion
Acknowledgments
The research work was supported by the Hellenic Foundation for Research and Innovation (HFRI) under the 3rd Call for HFRI PhD Fellowships (Fellowship Number 5537).
@@ A.2 Prompts
First, we will present the ZS prompt configuration used for generating the contextually reconstructed Question-Answer pairs.
@@
  Next, with minor modifications, we introduce the FS prompt configuration that follows the same logic but incorporates in-context examples to guide the model’s responses.
@@ B.2 RiddleSense
The first pipeline involves a more granular approach by splitting the contextual reconstructed question into subphrases based on punctuation or conjunctions. If this yields fewer than three distinct subphrases, we further split the sentence at the position of the word “and” to create additional segments. We also detect the presence of interrogative words, and if the question is purely descriptive without any direct question, we append the phrase “What am I?” at the end. This is not chosen arbitrarily but follows the standard structure of many riddles in this dataset, where "What am I?" is the common question leading to single-word answers. Now that we have broken the riddle into sub-phrases and appended the appropriate question, we can prompt the model to generate a wrong answer for each sub-phrase concatenated with the question. This method ensures that the distractors align with different parts of the riddle’s setting. However, since some sub-phrases may contain key ideas central to the riddle, the model may still generate answers too similar to the correct one. To address this issue, we incorporate an additional intermediate step utilizing zero-shot classification. Specifically, we use the facebook/bart-large-mnli model 555facebook/bart-large-mnli Lewis et al. (2019) provided by Hugging Face. This model classifies the correct answer into one of eight general categories: ’food’, ’person’, ’object’, ’animal’, ’nature’, ’time’, ’place’, ’concept’. These categories are chosen to be mutually exclusive to avoid overlap. In our approach, we first predict the category of the correct answer using this model. We then leverage this classification to guide the generation of distractors. For each sub-phrase concatenated with the question, we provide the LLM with the two most similar categories (excluding the correct answer’s category). The model is prompted with a system-user instruction to produce a correct answer of the given setting in the given category. This method ensures that the generated distractors are not only incorrect but also contextually aligned with the riddle, thus maintaining the challenge for the model. After applying filtering to ensure the distractors are distinct and relevant, we have produced several distractors. The LLMs utilized in this pipeline include the previously mentioned Llama3-8B AI@Meta (2024) and its larger counterpart, Llama3-70B AI@Meta (2024).